# Cross-Entropy Games for Language Models: From Implicit Knowledge to General Capability Measures

### v1

@@ 2.2.9. Two Simple Examples: Reverse Prompt and Paradox Game
Initialization: load a random $p$-token text into $s$ (e.g. generated by an LLM). Elicit $t$, a $q$-token move from the (unique) player. Give $-xent\left(s|t\right)$ to the player.
This first game is a simple (though typically hard) combinatorial optimization game about the implicit knowledge of $\mathcal{J}$: it essentially asks the player to find the best ‘summary’ (in a certain sense) from the point of view of the judge measure $\mathcal{S}_{\mathcal{J}}$. The reverse prompt game appears to be closely related to a number of interesting problems:
  • Compression problem: the question can be phrased as asking how to ‘pack’ as much information about $s$ as possible. An interesting variant is to take several texts $c_{1},\ldots,c_{n}$ and give $-\sum_{j=1}^{n}xent\left(c_{j}|t\right)$ to $black$.
@@
  • Problem solving: given a problem with an easy-to-check solution can one find a prompt (the solution) leading $\mathcal{J}$ to acknowledge that the solution is correct?
Generally speaking, the reverse-prompt game (as well as some variants of it) captures many features that we would like to access from an implicit measure point of view: surely a model can generate, given a prompt, but can it figure out what prompt could have made a certain output likely? Does it know what can cause it it to say something? As will be seen in Section 2.3, from this single game and two game axioms, one gets all the space of Xent Games.
@@
In practice, the game of Example 14 can admit, in its raw form, some unfortunately uninteresting solutions: for instance, repeating some low-probability tokens of the text $s$ can be a winning strategy, if $s$ contains rare enough word combinations. In Section 3.1, we discuss somehow less trivial variants that are more interesting to play (both for humans and models).
The second example is somehow simpler, though definitely less natural: it is a (and probably the simplest) paradoxical example that shows that shows a (theoretical) impossibility of models to have full access to the implicit measure. It is a zero-sum perfect information two-player game:
@@
This is a small twist on the paradox ‘the shortest sentence that cannot be described in ten words’ (that sentence cannot exist as such because, the quotemarked description would be itself a ten-word description of it).
@@
While interesting as a means to illustrate a point, the paradoxical game is an example of an ill-posed game (see 4.1.2 below), and as such it is not particularly interesting. Coupling it with some more constraints or rewards can make it well-posed and interesting, however (see Section 3.2 for instance).
@@ 2.3. Xent Space Characterization
Informally, the characterization result shows that the space of Xent Games is connected: by simple ‘moves’ in the game space, one can go from any Xent Game to any Xent Game.
@@ 2.3.1. Turn-By-Turn String Games: Definition and Composition
If one asks what is the natural space of games to be played for a fixed number of steps by a fixed number $n\geq 1$ of LLMs, given that all they do is to process text, one is naturally led to a space of games with the same structure as in Section 2.2.6, except that the rewards are general and that the string space operations are general. We call this space of general turn-by-turn $n$-player string games
@@
  • Elicits a move from a player, based on information made of a subset of the string space and on the player’s rewards so far, and of a subset of rewards of the other players.
@@
An important point about this (very large) space of games is that we can assume, without loss of generality that the set of players (i.e. their names) is the same for all games. Similarly, we can assume that the set of variables in string space is always the same (the variables are always initialized to the empty string by default). This leads to the fundamental property of composition of the operations (in a similar spirit as the so-called Open Games framework, [GHWZ16]): operations from two games can be composed sequentially.
@@
The question addressed in Section 2.3.3 is to characterize the Xent Games as being a relevant subspace of that of the general string games of Definition 19: it naturally follows from two stability properties.
@@ 2.3.2. Adversarial Rescaling and Zero-Summing Stability
From the composition property (Definition 20) follows the fact that rewards can be rescaled by an integer factor $\lambda\geq 0$: for $\lambda=0$, the reward step can be skipped, for $\lambda\geq 2$, one can put multiple copies of the reward step. If, inspired by the theory of Lagrange multipliers, we allow the reward rescaling factor to be chosen by an external adversarial agent conspiring against the player $P$ receiving the reward, which we call adversarial reward rescaling.
Like in the theory of Lagrange multiplier, an adversarial reward rescaling corresponds to transforming a reward step into a hard constraint that this reward must be nonnegative: if the reward is nonnegative, the adversary will simply multiply it by 00 (thus it becomes void), if the reward is negative, the adversary will multiply by $\lambda\to+\infty$, causing the player to lose. Again, like in the theory of Lagrange multipliers, the adversarial reward rescaling stability allows one to go from rewards to constraints, putting priorities on certain ‘rewards’ (or corresponding to ‘regulations’): first a mini-game must be won (against the ‘adversary’), and then if that mini-game is won, the string space is used to play another game. This leads us to the following formulation:
@@
A family $\mathcal{G}$ of string games is stable under (dynamic) adversarial (reward) rescaling if for each game $G\in\mathcal{G}$ and each reward step $R$ to a player $P$, the game $G^{\prime}$ where that reward $R$ is replaced by the constraint that the amount of $R$ must be nonnegative is still in $\mathcal{G}$.
@@
A family $\mathcal{G}$ of string games is stable under zero-summing if for each game $G\in\mathcal{G}$, the game $G_{0}$ where each player’s reward is followed by a reward of negative that amount to another player is still in $\mathcal{G}$.
@@ 2.3.3. Characterization Result
From the above three conditions (Definitions 20, 21, and 23) and the basic reverse prompt game (Example 14), we can identify the Xent Games (with no ensure fail allowed, see Remark 22), defined in Section 2.2:
@@
It is obvious from its design (in Section 2.2) that the space of Xent Games is stable under cat/cut updates, sequential composition, adversarial rescaling and zero-summing, and hence ΣXent  $\Sigma\subset\text{Xent\,\,Games}$. To show the reverse inclusion Xent  Games  $\text{Xent\,\,Games $\subset$ }\Sigma$ we must show that we can build any Xent Game from the properties of $\Sigma$. This is guaranteed by the from the following observations:
@@ 2.4. XGL: Xent Game Language
By their sequential design, Xent Games naturally lend themselves to a procedural description. In this subsection, we define a domain-specific language allowing one to represent the execution of a Xent Game as a program composed of simple instructions. This language, which we call XGL (Xent Game Language) is designed to maximize the simplicity of reading and definition of Xent Games, while at the same time maximizing the efficient automatic creation of such games (see Section 2.4.5).
@@ 2.4.1. XGL Design and Features
  • The string space is made of a fixed list of $32$ string registers grouped in types $s,t,x,y,p,c$, with each type coming with $4$ registers: the $4$ $s$-type registers are $s,s0,s1,s2$, the $4$ $t$-type registers are $t,t0,t1,t2$, etc. The $p$-type and $a,b,c$-type register have special rules:
@@
    • The players $black$, $white$ are a zero-sum pair (see reward comments below).
@@
  • The number of lines of an XGL game is capped to 64, which may however include looping for a fixed number of steps (these are hence not ‘truly’ conditionals, as they could be realized by a pre-processor), leading to a fixed maximal ‘unrolled’ length of 1024 steps:
@@
  • The complexity of the elicit calls, which include revealing specific elements of information to the elicited player and imposing ensure statements to the elicited is flattened:
@@
    • The constraints are enforced by ensure calls that follow the elicit; if the ensure condition fails, the execution jumps back to the last elicit call coming before the current ensure.
@@
    • As discussed above, $black$, $white$ are a zero-sum pair: any reward amount $\rho$ given to one player automatically comes with negative reward amount $-\rho$ to the other one; either player can be used for a one-player perfect information game (the rewards to the other are simply discarded, if it doesn’t play).
    • In addition to the values of the rewards being disclosed to the players, for each xent that is part of a reward, the corresponding atomic xents $axent$ are shared with the player if $s$ is $m$-token long and $t$ $n$-token long $axent\left(s,t\right)$ is the $m$-dimensional vector with the $i$-th coordinate of $axent\left(s,t\right)$ being defined as $-\log\mathbb{P}_{\mathcal{J}}\left\{X_{n+i}=s_{i}|t_{1},\ldots,t_{n},s_{1},\ldots,s_{i-1}\right\}$.
@@ 2.4.2. XGL Register and Instruction set
As discussed above, XGL is an assembly-like language, with each line containing an instruction, and being executed after the previous one. The instructions update the register and interact with the users. The registers and pre-defined variables set is the following
@@
  • $elicit$: updates string space from player’s moves. For instance $elicit(x,max\_len)$ takes the input of player $white$ (if not specified, the player is $black$) and stores the result into $x$, and $elicit(alice,x1,x2,x3,max\_len)$ asks for $3$ inputs $x1,x2,x3$ from $alice$ of lengths at most $max\_len$.
  • $ensure$: enforce conditions on the player’s moves (following an $elicit$). The $ensure$ calls take a list of booleans as argument, in particular those returned by $is\_true$ calls (see Section 2.4.3 below). For instance, $ensure(is\_true("num\_words<10",s))$ asks the model $\mathcal{J}$ to judge whether $s$ contains fewer than $10$ words.
  • $reward$: rewards a player (by default $white$) by taking $xent$-based signed sums (made of sums/differences of $xent,nex,xed,dex$, see Section 2.4.3 below) as inputs. For instance,
    • $reward(xent(s|t))$ gives $xent\left(s|t\right)$ to $white$ (the default player),
    • $reward(xed(s|t))$ gives $xed(s|t)=xent(s)-xent(s|t)$ to $white$,
@@
  • $assign$: updates string space (from string space itself, via cat/cut). For instance $assign\left(s=s1+s2\right)$ concatenates $s1$ and $s2$ and puts the result into $s$, and $assign\left(s0=c0/s\right)$ cuts $c0$ at the first occurrence of $s$ and puts what comes before into $s0$, while $assign(t0=c0//s)$ puts what comes after into $t0$.
  • $reveal$: string space registers to be shared with player (ahead of an $elicit$). For instance $reveal(alice,s2+t2)$ shares the string $s2+t2$ with Alice.
@@
  • $replay$: jumps to a previous flag. For instance, $replay(flag\_1,10)$ jumps to $flag\_1$ at most $10$ times.
@@ 2.4.3. Xent-based functions
  • We have $xent\left(s|t,o\right)=-\log\mathbb{P}_{\mathcal{J},o}\left\{s|t\right\}$ where $o$ is a general set of customization pre-instructions for $\mathcal{J}$ that the model is free to interpret (and that will typically come before the pre-prompt $t$), which must be an in-line constant. By default, if $t$ is empty $xent\left(s,o\right)=-\log\mathbb{P}_{\mathcal{J},o}\left\{s\right\}$; if $s$ is empty, $xent\left(s|t,o\right)$ is zero.
@@
The $ensure$ calls take booleans which can be the results of explicit xent comparisons (e.g. $ensure(xent(s|t)<xent(s))$) or implicit ones, via built-in $is\_true$ and $is\_false$ calls:
@@ 2.4.4. XGL Metadata and Pre-Game Code
Most of the XGL metadata consists of data to link $\mathcal{J}$ to a specific LLM, and to link the players to various LLM agents, as discussed Xent Game description above (Section 2.2.1). This data is set in a dictionary (if modifications to default values are needed) and is typically expected to be fixed once and for all for the Xent space, and it includes the prompts to fill the a, b, c registers. The only part of the metadata that is absent is the session seed, which is set externally. Like the code, the metadata is shared with the all the players.
The only difference with the list of Section 2.2.1 is that much of the metadata described there appears in-line in the game code via the use of quote-marked string constants (this choice is to made for improved readability by the players).
One important design element is that the metadata can be sequentially defined by several dictionaries, where each dictionary can override the variables defined by the dictionaries that came before (in the order presented below):
@@ 2.4.5. XGL as a Target Language for LLMs
In addition to readability and ease of execution, the XGL language is designed from the beginning to facilitate the automatic generation of games, which are instrumental in Section 6.3 below to discover new games relevant to measuring general model abilities, i.e. to be a good target language for LLM generation. The structure of the language is designed to help with the spontaneous discovery of new games by abiding to the following principles:
  • An XGL program is valid if and only if all its lines are individually valid XGL lines: this allows one to e.g. ‘cross-breed’ games.
  • The structure of the $reward$ and $ensure$ are similar, and they take as inputs xents.
  • All the variables are named in advance, and they are string only.

### v2

@@ 2.2.9. Two Simple Examples: Reverse Prompt and Paradox Game
Initialization: load a random (e.g. generated by an LLM) $p$-token text into $s$. Elicit $t$, a $q$-token move from the (unique) player. Reward $-xent\left(s|t\right)$ to the player.
This first game is a simple (though typically hard) combinatorial optimization game about the implicit knowledge of $\mathcal{J}$: it essentially asks the player to find the best ‘summary’ of $s$ from the point of view of the judge measure $\mathcal{S}_{\mathcal{J}}$. The reverse prompt game appears to be closely related to a number of interesting problems:
  • Compression problem: the question can be phrased as asking how to ‘pack’ as much information as possible about $s$. An interesting variant is to take several texts $c_{1},\ldots,c_{n}$ and give $-\sum_{j=1}^{n}xent\left(c_{j}|t\right)$ to the player.
@@
  • Problem solving: given a problem with an easy-to-check solution, can one find a prompt (the solution) which leads $\mathcal{J}$ to acknowledge that the solution is correct?
Generally speaking, the reverse-prompt game (as well as some variants of it) captures many features that we would like to access from an implicit measure point of view: surely a model can generate, given a prompt, but can it figure out what prompt could have made a certain output likely? Does it know what can cause it it to say something? As will be seen in Section 2.3, from this single game and two game axioms, one gets the entire space of Xent Games.
@@
In practice, the game of Example 14 can admit, in its raw form, some unfortunately uninteresting solutions. For instance, repeating some low-probability tokens of the text $s$ can be a winning strategy, if $s$ contains rare enough word combinations. In Section 3.1, we discuss some less trivial variants that are more interesting to play (both for humans and models).
The second example is somehow simpler, though definitely less natural: it is a paradoxical example (and probably the simplest one) that shows that shows the (theoretical) impossibility of models having full access to the implicit measure. It is a zero-sum perfect information two-player game:
@@
This is a small twist on the paradox ‘the shortest sentence that cannot be described in ten words’ (that sentence cannot exist as such because the quotemarked description would be itself a ten-word description of it).
@@
While interesting as a means to illustrate a point, the paradoxical game is an example of an ill-posed game (see 4.1.2 below), and as such it is not particularly interesting. However, coupling it with some more constraints or rewards can make it well-posed and interesting (see Section 3.2 for instance).
@@ 2.3. Xent Space Characterization
Informally, the characterization result shows that the space of Xent Games is connected: by simple ‘moves’ in the game space, one can go from any Xent Game to any other Xent Game.
@@ 2.3.1. Turn-By-Turn String Games: Definition and Composition
If one asks what is the natural space of games to be played for a fixed number of steps by a fixed number $n\geq 1$ of LLMs, given that all they do is to process text, one is naturally led to a space of games with the same structure as in Section 2.2.6, except that the rewards are general and that the string space operations are general.
@@
  • Elicits a move from a player, based on information made up of a subset of the string space, the player’s rewards so far, and a subset of rewards of the other players.
@@
An important point about this (very large) space of games is that we can assume, without loss of generality that the set of players (i.e. their names) is the same for all games. Similarly, we can assume that the set of variables in the string space is always the same (the variables are always initialized to the empty string by default). This leads to the fundamental property of composition of operations (in a similar spirit as the so-called Open Games framework, [GHWZ16]): operations from two games can be composed sequentially.
@@
The question addressed in Section 2.3.3 looks to characterize Xent Games as being a relevant subspace of the general string games of Definition 19. This naturally follows from two stability properties:
@@ 2.3.2. Adversarial Rescaling and Zero-Summing Stability
From the composition property (Definition 20) follows the fact that rewards can be rescaled by an integer factor $\lambda\geq 0$: for $\lambda=0$, the reward step can be skipped, for $\lambda\geq 2$, one can put multiple copies of the reward step. Inspired by the theory of Lagrange multipliers, we allow the reward rescaling factor to be chosen by an external adversarial agent conspiring against the player $P$ receiving the reward, which we call adversarial reward rescaling.
Like in the theory of Lagrange multiplier, an adversarial reward rescaling corresponds to transforming a reward step into a hard constraint that this reward must be nonnegative: if the reward is nonnegative, the adversary will simply multiply it by 00 (thus it becomes void), if the reward is negative, the adversary will multiply by $\lambda\to+\infty$, causing the player to lose. Again, like in the theory of Lagrange multipliers, the adversarial reward rescaling stability allows one to go from rewards to constraints, putting priorities on certain ‘rewards’ (corresponding to ‘regulations’): first a mini-game must be won (against the ‘adversary’), and then, if that mini-game is won, the string space is used to play another game. This leads us to the following formulation:
@@
A family $\mathcal{G}$ of string games is stable under (dynamic) adversarial (reward) rescaling if for each game $G\in\mathcal{G}$ and each reward $R$ to a player $P$, the game $G^{\prime}$ where that reward $R$ is replaced by the constraint that the amount of $R$ must be nonnegative, is still in $\mathcal{G}$.
@@
A family $\mathcal{G}$ of string games is stable under zero-summing if, for each game $G\in\mathcal{G}$, the game $G_{0}$, where each player’s reward is followed by a reward of negative that amount to another player, is still in $\mathcal{G}$.
@@ 2.3.3. Characterization Result
From the above three conditions (Definitions 20, 21, and 23) and the basic reverse prompt game (Example 14), we can identify the Xent Game space (with no ensure fail allowed, see Remark 22), defined in Section 2.2:
@@
It is obvious from its design (in Section 2.2) that the space of Xent Games is stable under cat/cut updates, sequential composition, adversarial rescaling and zero-summing, and hence ΣXent  $\Sigma\subset\text{Xent\,\,Games}$. To show the reverse inclusion Xent  Games  $\text{Xent\,\,Games $\subset$ }\Sigma$ we must show that we can build any Xent Game from the properties of $\Sigma$. This is guaranteed by following observations:
@@ 2.4. XGL: Xent Game Language
By their sequential design, Xent Games naturally lend themselves to a procedural description. In this subsection, we define a domain-specific language allowing one to represent a Xent Game as a program composed of simple instructions. This language, which we call XGL (Xent Game Language) is designed to maximize the simplicity of reading and defining Xent Games, while at the same time maximizing the efficient automatic creation of such games (see Section 2.4.5). An implementation of XGL is available in the xega repository on GitHub.
@@ 2.4.1. XGL Design and Features
  • The string space is made of a fixed list of $32$ string registers grouped in types $a,b,s,t,x,y,p,c$, with each type coming with $4$ registers: the $4$ $s$-type registers are $s,s0,s1,s2$, the $4$ $t$-type registers are $t,t0,t1,t2$, etc. The $p$-type and $a,b,c$-type register have special rules:
@@
    • The players $black$ and $white$ are a zero-sum pair (see reward comments below).
@@
  • The number of lines of an XGL game is capped to 64, which may however include looping for a fixed number of steps (these are hence not ‘truly’ conditionals, as they could be realized by a pre-processor), leading to a fixed maximal ‘unrolled’ length of 1024 steps. Looping occurs:
@@
  • The complexity of the elicit calls, which include revealing specific elements of information to the elicited player and imposing ensure statements to the elicited response, is flattened:
@@
    • The constraints are enforced by ensure calls that follow an elicit; if the ensure condition fails, execution jumps back to the last elicit call coming before the current ensure.
@@
    • As discussed above, $black$ and $white$ are a zero-sum pair: any reward amount $\rho$ given to one player automatically comes with negative reward amount $-\rho$ to the other one; either player can be used for a one-player perfect information game (the rewards to the other are simply discarded if they don’t play).
    • In addition to the values of the rewards being disclosed to the players, for each xent that is part of a reward, the corresponding atomic xents $axent$ are shared with the player. If $s$ is $m$-token long and $t$ $n$-token long, then $axent\left(s,t\right)$ is the $m$-dimensional vector with the $i$-th coordinate of $axent\left(s,t\right)$ being defined as $-\log\mathbb{P}_{\mathcal{J}}\left\{X_{n+i}=s_{i}|t_{1},\ldots,t_{n},s_{1},\ldots,s_{i-1}\right\}$.
@@ 2.4.2. XGL Register and Instruction set
As discussed above, XGL is an assembly-like language. Lines are executed sequentially (top-to-bottom), with each line containing a single instruction. XGL instructions update the registers and interact with the players. The registers and pre-defined variables set are the following:
@@
  • $elicit$: updates string space from player’s moves. For instance $elicit(x,20)$ asks for input of max length 20 from player $black$ (if not specified, the player is $black$) and stores the result into $x$, and $elicit(alice,x1,x2,x3,10)$ asks for $3$ inputs $x1,x2,x3$ from $alice$ of lengths at most 10.
  • $ensure$: enforce conditions on the player’s moves (following an $elicit$). The $ensure$ calls take a list of boolean expressions as arguments, in particular those returned by $is\_true$ calls (see Section 2.4.3 below). For instance, $ensure(is\_true("num\_words<10",s))$ asks the model $\mathcal{J}$ to determine whether $s$ contains fewer than $10$ words.
  • $reward$: rewards a player (by default $black$) by taking $xent$-based signed sums (made of sums/differences of $xent,nex,xed,dex$, see Section 2.4.3 below) as inputs. For instance,
    • $reward(xent(s|t))$ gives $xent\left(s|t\right)$ to $black$ (the default player),
    • $reward(xed(s|t))$ gives $xed(s|t)=xent(s)-xent(s|t)$ to $black$,
@@
  • $assign$: updates string space (from string space itself, via cat/cut). For instance $assign\left(s=s1+s2\right)$ concatenates $s1$ and $s2$ and puts the result into $s$, and $assign\left(s0=c0//s\right)$ cuts $c0$ at the first occurrence of $s$ and puts what comes before into $s0$, while $assign(t0=c0\%s)$ puts what comes after into $t0$.
  • $reveal$: marks string space registers to be shared with player (ahead of an $elicit$). For instance $reveal(alice,s2+t2)$ shares the string $s2+t2$ with Alice.
@@
  • $replay$: jumps execution to a previous flag. For instance, $replay(flag\_1,10)$ jumps to $flag\_1$ at most $10$ times.
@@ 2.4.3. Xent-based functions
  • We have $xent\left(s|t,o\right)=-\log\mathbb{P}_{\mathcal{J},o}\left\{s|t\right\}$ where $o$ is a general set of customizable pre-instructions for $\mathcal{J}$ that that will come before the pre-prompt $t$. Note that $o$ must be an in-line constant. By default, if $t$ is empty $xent\left(s,o\right)=-\log\mathbb{P}_{\mathcal{J},o}\left\{s\right\}$; if $s$ is empty, $xent\left(s|t,o\right)$ is zero.
@@
The $ensure$ calls take booleans which can be the result of explicit xent comparisons (e.g. $ensure(xent(s|t)<xent(s))$) or implicit ones, via built-in $is\_true$ and $is\_false$ calls:
@@ 2.4.4. XGL Metadata and Pre-Game Code
Most of the XGL metadata consists of data to link $\mathcal{J}$ to a specific LLM, and to link the players to various LLM agents, as discussed Xent Game description above (Section 2.2.1). This data is set in a dictionary (if modifications to default values are needed) and it includes the prompts to fill the a, b, c registers. The only part of the metadata that is absent is the session seed, which is set externally. Like the code, the metadata is shared with all players.
The only difference with the list of Section 2.2.1 is that much of the metadata described there appears in-line in the game code via the use of quote-marked string constants. This choice is to made to improve readability for the players.
One important design element is that the metadata can be sequentially defined by several dictionaries, where each dictionary can override the variables defined by the dictionaries that came before in the order presented below:
@@ 2.4.5. XGL as a Target Language for LLMs
In addition to readability and ease of execution, XGL is designed from the beginning to facilitate the automatic generation of games, which is instrumental in Section 6.3 below to discover new games relevant to measuring general model abilities. Put simply, XGL is designed to be a good target language for LLM generation. The structure of the language is designed to help with the spontaneous discovery of new games by abiding by the following principles:
  • An XGL program is valid if and only if all its lines are individually valid XGL lines. This allows one to e.g. ‘cross-breed’ games.
  • The structure of the $reward$ and $ensure$ are similar and they both take xents as inputs.
  • All the variables are named in advance and they are string typed only.